Calculate the first-order (n=1) Bragg diffraction angle for 117.0 pm X-rays passing through a crystalline lattice having a spacing between adjacent molecular layers of 178.0 pm.
1
Expert's answer
2019-09-16T04:02:50-0400

Bragg's Law refers to the simple equation:

"n\lambda=2d\sin(\theta)\\here,\\n=1\\\lambda=117\ pm\\d=space\ bw\ adjacent\ molecular\ layer=178\ pm\\\sin(\theta)=\frac{1\times117}{2\times178}=0.32865\\"

So,

"\theta=19.1868\degree"


first-order (n=1) Bragg diffraction angle = "19.1868\degree"